# Multi-Level Querying using A Knowledge Pyramid

### v2

@@ 5.3 Comparison to SOTA Methods
We compare PolyRAG with five groups of state-of-the-art (SOTA) methods, including 1) Frozen-LLMs that are pretrained LLMs of frozen parameters and no external knowledge has been used for querying. 2) SFT (Supervised Fine-Tuning) that are LLMs that undergo supervised fine-tuning on domain-specific datasets. We include two latest implementation of Finance-LLM and Finance-Chat Cheng et al. (2024), both of which are trained on financial corpora. Additionally, we explore variants of SFT that employ LoRA Hu et al. (2022) instead of the entire models. 3) NaiveRAG proposed in Lewis et al. (2020a). 4) Advanced RAG that include two recent models, namely ColBERTv2 Santhanam et al. (2022) and Bge-reranker Xiao et al. (2023). 5) KG-Augmented LLMs that consists of models proposed in Baek et al. (2023b) that integrate knowledge graphs (KGs) into the LLMs. Various LLMs have been explored as the backbones, including LLaMA Touvron et al. (2023a), LLaMA2 Touvron et al. (2023b), Vicuna Chiang et al. (2023), Gemini Team et al. (2023), GPT-3.5 and GPT-4. By combining them with the five groups and PolyRAG, this results in a comprehensive set of 23 methods for performance comparison.
The results are shown in Table 1. It is clear that PolyRAG surpasses other SOTA methods across the five groups, different backbones, and both benchmarks. An encouraging finding is that PolyRAG demonstrates a superior balance between precision and recall. Among the improvements of PolyRAG over others, the gain on the precision ($22.0\%\pm 17.3$) is more pronounced than that of recall ($18.4\%\pm 18.6$). With the incorporation of PolyRAG, GPT-4 (GPT-3.5) has achieved the highest gains by $338.6\%$ ($133.2\%$) over its frozen model and $73.1\%$ ($69.8\%$) over its NaiveRAG run on the AcadChall benchmark. The improvement may result from the fact that around 43.7% (38.4%) of questions are addressed at the Ontology layer, and around 42.3% (44.8%) are at the KG layer on the AcadChall (R-FLUE-FiQA) benchmark.
5.4 Knowledge Completion and Condensation
We have conducted experiments to evaluate the effectiveness of knowledge completion (CPL) and condensation (CND) by integrating them with a baseline model (PolyRAG without these two modules). The results are presented in Table 2. Please note that due to space limitations, we only list results for four specific backbones. For a comprehensive performance analysis with Vicuna-13B and LLaMA2-70B backbones, please refer to the Appendix C. The results clearly demonstrate the effectiveness of both completion and condensation in enhancing performance, whether used individually (CPL: $3.8\%\pm 6.1$ performance gain, CND: $2.2\%\pm 3.3$ performance gain) or in combination ($6.5\%\pm 6.4$ performance gain over the baseline). Our statistics indicate that with completion, queries resolved at the Ontology layer is increased by 5%. With condensation, more than 27% of queries are resolved at the KG layer.
5.5 Influence of Knowledge Layers
We have examined the influence of different knowledge layers on various backbones. The results are presented in Table 3. It is obvious that the Ontology layer plays a prominent role as a knowledge base in enhancing precision, as evidenced by a precision gain of $25.8\%\pm 7.4$ when combined with the raw text layer and $13.7\%\pm 10.3$ when combined with the KG layer. On the other hand, the KG layer demonstrates a balanced impact on improving either precision or recall. These observations align with our early discussions.
6 Conclusion
This research has described a preliminary investigation that provides a multi-level query approach to domain-specific question answering. In this context, we addressed two significant challenges: the problem of prioritizing knowledge in a certain area and the presence of noise in retrieval situations. In order to overcome the limitations, we introduce Knowledge Pyramid multi-level retrieval framework as PolyRAG. PolyRAG utilizes a sequential retrieval approach and prioritizes knowledge extraction. This approach is designed to tackle the challenges posed by the high demand for dense knowledge in domain-specific scenarios and the distractions caused by noisy contexts. The accuracy of PolyRAG has been validated through extensive tests done in AcadChall and R-FLUE-FiQA, surpassing earlier approaches and generating state-of-the-art results.
Limitations
One limitation to consider is that Language Models (LLMs) do not strictly adhere to the SparQL syntax, which can result in typographical errors in the queries. Further investigation into this aspect is warranted.
Ethics Statement
This research paper adheres to ethical considerations throughout its methodology and findings. The study primarily focuses on the development and evaluation of a retrieval-augmented generation framework, PolyRAG, for domain-specific knowledge retrieval. The research involves the use of existing datasets and benchmarks, as well as the creation of two new benchmarks in the academic and financial domains.
The authors ensure proper citation and acknowledgment of the data sources to maintain transparency and intellectual property rights. In terms of human subjects, this research does not involve any direct human participation, personal data collection, or sensitive information. Therefore, ethical concerns related to informed consent, privacy, and confidentiality are not applicable in this context.
Appendix
Appendix A Knowledge Graph Layer Construction
This section provides the detailed implantations of the four functions for constructing the knowledge graph layer in Section 3.1.
Appendix B AcadChall and R-FLUE-FiQA Benchmarks
Table 4 lists the data statistics of two benchmarks. We use Vicuna-13B Chiang et al. (2023) to build the knowledge pyramid for two datasets according to the methods illustrated in Section 3.1. Specifically, the AcadChall dataset is created by crawling public information from the official websites of various departments at XXX University. It encompasses comprehensive information about 1,319 staff members, 2,061 courses, 31 departments, classrooms, and facilities, with over 5k corpus entries. Taking the staff member category as an example, it includes not only personal information, contact details, positions held at the university, and department affiliations but also specific academic knowledge, such as staff research interests, alma maters, degrees, research experience, and published papers, which is hard to be captured by general LLMs. In addition, we manually create 512 multi-choice QA pairs. Each question corresponds to eight options, and the answer could be one or more options. For example, for the question “Which staff member in the Art department graduated from Cambridge University?”, there could be multiple staff in the provided options that meet the information need. For the R-FLUE-FiQA dataset, we adopt the test set from the FLUE benchmark Shah et al. (2022), which consists of 1,705 open-ended long-form QA pairs.
Table 4: Dataset statistics of the two benchmarks.
Dataset AcadChall R-FLUE-FiQA
Domain Academia Finance
#Raw texts 5,019 17,072
Ontologies
   #Instances 25,481 225,085
#KG Triples 27,920 68,737
#QA pairs 512 1,705
QA type MCQ&MAQ Open-ended
Table 5: The impact of Knowledge Completion (CPL) and Condensation (CND) by precison, recall, and F1 scores.
Method Vicuna-7b Vicuna-13b LLaMA2-70b Gemini-Pro GPT-3.5 GPT4
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Naive RAG 0.3621 0.8386 0.4673 0.3665 0.9043 0.4923 0.5673 0.3893 0.4193 0.5152 0.3968 0.4214 0.5105 0.5791 0.5315 0.5032 0.5094 0.4651
Baseline 0.5306 0.8337 0.5734 0.4230 0.8848 0.5309 0.7661 0.6479 0.6678 0.7956 0.6513 0.6899 0.7855 0.7806 0.7905 0.7661 0.8206 0.7639
+ CPL 0.6888 0.8325 0.5874 0.4289 0.8045 0.5595 0.8633 0.6836 0.7267 0.8217 0.8597 0.7869 0.8044 0.8152 0.7939 0.8063 0.7909 0.7763
+ CND 0.6557 0.7599 0.5766 0.4236 0.8075 0.5557 0.8595 0.6445 0.7166 0.8036 0.8737 0.7439 0.7985 0.8027 0.7943 0.8435 0.7853 0.7873
+ CPL + CND 0.5306 0.8337 0.5967 0.4330 0.8148 0.5655 0.8838 0.6821 0.7380 0.8372 0.8419 0.8107 0.8667 0.7860 0.8039 0.8711 0.7992 0.8109
Table 6: The influence of different knowledge layers on the AcadChall benchmark.
Data Source Vicuna-7b Vicuna-13b LLaMA2-70b Gemini-Pro GPT-3.5 GPT4
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Ontology 0.4383 0.6242 0.4445 0.3363 0.5811 0.3818 0.5549 0.4501 0.4501 0.3386 0.3089 0.3157 0.4107 0.3945 0.3979 0.3500 0.4369 0.3677
KG 0.4365 0.8581 0.5133 0.3446 0.8945 0.4742 0.6666 0.6490 0.6080 0.6388 0.4591 0.5099 0.7326 0.6340 0.6433 0.6327 0.7065 0.6238
Raw Texts 0.3621 0.8386 0.4673 0.3665 0.9043 0.4923 0.5673 0.3893 0.4193 0.5152 0.3968 0.4214 0.5105 0.5791 0.5315 0.5032 0.5094 0.4651
Ontology + Raw Texts 0.4535 0.8393 0.5368 0.4100 0.8908 0.5187 0.6198 0.5244 0.5205 0.5905 0.4662 0.5013 0.6890 0.5818 0.6020 0.6464 0.6779 0.6303
KG + Raw Texts 0.3673 0.8758 0.4840 0.3747 0.9050 0.4919 0.7421 0.5083 0.5505 0.6890 0.5428 0.5821 0.8339 0.6622 0.7071 0.8187 0.7729 0.7636
Ontology + KG 0.5251 0.8266 0.5608 0.4081 0.8814 0.5166 0.6833 0.5980 0.5922 0.7755 0.6697 0.6709 0.8567 0.7710 0.7939 0.6083 0.9032 0.6871
PolyRAG 0.5306 0.8337 0.5967 0.4330 0.8148 0.5655 0.8838 0.6821 0.7380 0.8372 0.8419 0.8107 0.8667 0.7860 0.8039 0.8711 0.7992 0.8109
Appendix C Discussion on Knowledge Completion and Condensation
@@
Table 5 verifies the effectiveness of our proposed knowledge completion and condensation techniques over multiple LLM backends. Specifically, the multi-level querying on the knowledge pyramid without completion and condensation techniques is used as the baseline. In the baseline, around 30% of questions are answered by more structured data (i.e., Ontology or KG). After applying the completion and condensation techniques, the percentage increases to 64%. For example, the relation “has published in”, which frequently appears in the KG layer, has been used to complete the Ontology layer. It has resulted in those questions related to “publications” being answered by the Ontology layer. Similarly, with knowledge condensation, an increase of around 27% of questions are answered by the KG layer. For example, as “has published research in”, “has published research articles on”, “has published a paper in”, “has work published in” has been condensed tohas published research in ” or “has published research articles in”, the number of tokens that sent to LLM for QA is decreased. Having a compact prompt has eased a smaller LLM backend, such as Vicuna-7B, in understanding the context, and we found that it could lead to a better question-answering result. Overall, with the completion/condensation technique, precision is consistently improved over the baseline across different LLM backends. With the collaboration of completion and condensation techniques, the performance boost is more significant.
Appendix D Influence of knowledge layers over backbones
Table 6 lists the impact of different knowledge layers on more LLM backbones than the results presented in Section 5.5. The proposed knowledge pyramid significantly improves the precision of the raw text layer across different backbones. With more powerful backbones, such as Genmini-Pro, GPT-3.5, and GPT-4, the performance gains are larger (i.e., over 0.3). Besides, higher precision could be obtained when combining the Ontology layer with the KG or raw texts layer. A similar observation happens when combing the KG layer with raw text or the Ontology layer across different backbones. These are consistent with the observations in Section 5.5.

### v3

@@ V-C Comparison to SOTA Methods
We compare PolyRAG with five groups of SOTA methods, including 1) Frozen-LLMs that are pretrained LLMs with frozen parameters and no external context knowledge. 2) SFT (Supervised Fine-Tuning) that are LLMs undergoing supervised fine-tuning on domain-specific datasets. We include two latest implementations of Finance-LLM and Finance-Chat [58], both of which are trained on financial corpora. Additionally, we explore LoRA-based [59] SFT. 3) NaiveRAG proposed in [21]. 4) Advanced RAG that includes two recent models, namely ColBERTv2 [60] and Bge-reranker [38]. 5) KG-Augmented LLMs that consist of models proposed in [45] that integrate knowledge graphs (KGs) into the LLMs. We also explored different LLM backbones, including LLaMA [2], LLaMA2 [61], Vicuna [62], Gemini [63], GPT-3.5 and GPT-4. By combining them with the five groups and PolyRAG, we conduct a comprehensive set of 23 methods for performance comparison.
The results of our experiments are presented in Table II, which showcases the performance comparison between PolyRAG and other state-of-the-art (SOTA) methods across five different groups, various backbone models, and both benchmarks: AcadChall and R-FLUE-FiQA. From the table, it is evident that PolyRAG consistently outperforms other methods in terms of key evaluation metrics such as precision, recall, and F1 scores across all configurations. An encouraging finding from the results is that PolyRAG demonstrates a superior balance between precision and recall compared to other methods. Specifically, among the improvements of PolyRAG over others, the gain in precision ($22.0\%\pm 17.3\%$) is more pronounced than that of recall ($18.4\%\pm 18.6\%$). This indicates that PolyRAG is particularly effective at accurately retrieving relevant information while minimizing false positives, which is crucial for applications requiring high precision.
Furthermore, with the incorporation of PolyRAG, advanced language models like GPT-4 and GPT-3.5 have achieved significant performance gains. Notably, GPT-4 has achieved a remarkable gain of $338.6\%$ over its frozen model and $73.1\%$ over its NaiveRAG run on the AcadChall benchmark. Similarly, GPT-3.5 has achieved gains of $133.2\%$ over its frozen model and $69.8\%$ over its NaiveRAG run on the same benchmark.
The notable performance improvements may result from PolyRAG’s effective utilization of the hierarchical knowledge pyramid, particularly the higher layers. Our analysis indicates that around $43.7\%$ ($38.4\%$) of questions are addressed at the Ontology layer, and approximately $42.3\%$ ($44.8\%$) are resolved at the KG layer on the AcadChall (R-FLUE-FiQA) benchmark. This suggests that a significant portion of queries benefit from the structured and semantically rich information provided by the Ontology and KG layers. By efficiently leveraging these layers, PolyRAG enhances both precision and recall, leading to significant overall performance gains compared to other methods.
TABLE III: Ablation results on the AcadChall shows the impact of: (1) Completion (CPL) and Condensation (CDN); (2) knowledge layers, including various combinations of Ontology (Ont.), Knowledge Graph (KG), and Raw Text (Raw.).
Model Metric Knowledge CPL and CND Influence of Knowledge Layers
Naive Baseline +CPL +CND +CPL Ont. KG Raw. Ont. KG Ont. PolyRAG
RAG +CND +Raw. +Raw. +KG
Vicuna-7b Precision 0.3621 0.5306 0.6888 0.6557 0.5306 0.4383 0.4365 0.3621 0.4535 0.3673 0.5251 0.5306
Recall 0.8386 0.8337 0.8325 0.7599 0.8337 0.6242 0.8581 0.8386 0.8393 0.8758 0.8266 0.8337
F1 0.4673 0.5734 0.5874 0.5766 0.5967 0.4445 0.5133 0.4673 0.5368 0.4840 0.5608 0.5967
Vicuna-13b Precision 0.3665 0.4230 0.4289 0.4236 0.4330 0.3363 0.3446 0.3665 0.4100 0.3747 0.4081 0.4330
Recall 0.9043 0.8848 0.8045 0.8075 0.8148 0.5811 0.8945 0.9043 0.8908 0.9050 0.8814 0.8148
F1 0.4923 0.5309 0.5595 0.5557 0.5655 0.3818 0.4742 0.4923 0.5187 0.4919 0.5166 0.5655
LLaMA2-70b Precision 0.5673 0.7661 0.8633 0.8595 0.8838 0.5549 0.6666 0.5673 0.6198 0.7421 0.6833 0.8838
Recall 0.3893 0.6479 0.6836 0.6445 0.6821 0.4501 0.6490 0.3893 0.5244 0.5083 0.5980 0.6821
F1 0.4193 0.6678 0.7267 0.7166 0.7380 0.4501 0.6080 0.4193 0.5205 0.5505 0.5922 0.7380
Gemini-Pro Precision 0.5152 0.7956 0.8217 0.8036 0.8372 0.3386 0.6388 0.5152 0.5905 0.6890 0.7755 0.8372
Recall 0.3968 0.6513 0.8597 0.8737 0.8419 0.3089 0.4591 0.3968 0.4662 0.5428 0.6697 0.8419
F1 0.4214 0.6899 0.7869 0.7439 0.8107 0.3157 0.5099 0.4214 0.5013 0.5821 0.6709 0.8107
GPT-3.5 Precision 0.5105 0.7855 0.8044 0.7985 0.8667 0.4107 0.7326 0.5105 0.6890 0.8339 0.8567 0.8667
Recall 0.5791 0.7806 0.8152 0.8027 0.7860 0.3945 0.6340 0.5791 0.5818 0.6622 0.7710 0.7860
F1 0.5315 0.7905 0.7939 0.7943 0.8039 0.3979 0.6433 0.5315 0.6020 0.7071 0.7939 0.8039
GPT-4 Precision 0.5032 0.7661 0.8063 0.8435 0.8711 0.3500 0.6327 0.5032 0.6464 0.8187 0.6083 0.8711
Recall 0.5094 0.8206 0.7909 0.7853 0.7992 0.4369 0.7065 0.5094 0.6779 0.7729 0.9032 0.7992
F1 0.4651 0.7639 0.7763 0.7873 0.8109 0.3677 0.6238 0.4651 0.6303 0.7636 0.6871 0.8109
V-D Knowledge Completion and Condensation
We have conducted extensive experiments to evaluate the effectiveness of our proposed Knowledge Completion (CPL) and Knowledge Condensation (CND) techniques by integrating them with a baseline model (PolyRAG without these two modules). The results are presented in the left part of Table III, which verifies the effectiveness of our methods across multiple LLM backends, including Vicuna-7b, Vicuna-13b, LLaMA2-70b, Gemini-Pro, GPT-3.5, and GPT-4. Specifically, the baseline model utilizes multi-level querying on the knowledge pyramid without the completion and condensation techniques.
The data in Table III clearly demonstrate that both completion and condensation significantly enhance performance. When used individually, CPL yields an average F1 score improvement of $3.8\%\pm 6.1\%$ over the baseline, while CND results in a gain of $2.2\%\pm 3.3\%$. Notably, when both CPL and CND are combined, the performance improvement increases to $6.5\%\pm 6.4\%$, showcasing a more significant boost due to their collaborative effect. Precision is consistently improved across different LLM backends, highlighting the effectiveness of our approach.
@@ V-D Knowledge Completion and Condensation
In the baseline, around 30% of questions are answered using more structured data (i.e., Ontology or KG layers). After applying the completion and condensation techniques, this percentage increases significantly to 64%. For example, with Knowledge Completion, the relation “has published in”, which frequently appears in the KG layer, has been integrated into the Ontology layer. This integration results in questions related to “publications” being answered more effectively by the Ontology layer, improving both precision and recall. Similarly, with Knowledge Condensation, there is an increase of around 27% of questions being resolved at the KG layer. By condensing similar relations such as “has published research in”, “has published research articles on”, “has published a paper in”, and “has work published in” into unified representations like “has published research in” or “has published research articles in”, we reduce the number of tokens sent to the LLM for QA. This reduction not only streamlines the context provided to the LLM but also aids smaller LLM backends, such as Vicuna-7B, in understanding the context more effectively, leading to better question-answering results.
Overall, the collaborative use of both techniques results in more significant performance boosts, as evidenced by the highest F1 scores achieved when both CPL and CND are applied together. For instance, in the case of GPT-4, the F1 score increases from 0.7639 (baseline) to 0.8109 with both CPL and CND. These results underscore the complementary benefits of integrating both methods, enhancing both the richness and compactness of the knowledge representation, and ultimately leading to more accurate and efficient question-answering performance across different language models.
V-E Influence of Knowledge Layers
We have conducted a comprehensive examination of the influence of different knowledge layers on various language model backbones. The results, presented in the right part of Table III, demonstrate the significant impact that integrating different layers of knowledge has on the performance of language models.
It is evident from the results that the Ontology layer plays a prominent role in enhancing precision across different backbones. For instance, when the Ontology layer is combined with the raw text layer, there is a substantial average precision gain of $25.8\%\pm 7.4\%$. Specifically, for Vicuna-7b, the precision increases from 36.21% (using raw texts alone) to 45.35% when combined with the Ontology layer. Similarly, for GPT-3.5, the precision improves from 51.05% to 68.90%, and for GPT-4, from 50.32% to 64.64%. When the Ontology layer is combined with the KG layer, we observe an average precision gain of $13.7\%\pm 10.3\%$, indicating that the Ontology layer enhances the structured knowledge provided by the KG layer, leading to improved precision.
On the other hand, the KG layer demonstrates a balanced impact on improving both precision and recall. The KG layer alone achieves high recall rates, such as 64.9% for LLaMA2-70b. This indicates that the KG layer effectively enhances the models’ ability to retrieve relevant information due to its rich semantic relationships. Additionally, the KG layer contributes to reasonable precision improvements. For example, for LLaMA2-70b, the precision with the KG layer is 66.66%, which is higher than using raw texts (56.73%) or the Ontology layer (55.49%). Notably, Vicuna-7b and Vicuna-13b, which are among the highest recall scores across all data sources and backbones, tend to select all the choices during the MCQ and MAQ questions and thus lead to an unreal high recall. This analysis verify our claim that balancing precision and recall brings significant meaning to these tasks.
These observations align with our earlier discussions about the benefits of leveraging multiple knowledge layers. The proposed knowledge pyramid significantly improves language model performance by utilizing each layer’s complementary strengths. The Ontology layer enhances precision with its structured and precise definitions, while the KG layer contributes to improved recall through its rich semantic connections. The models achieve better overall performance when combined with raw texts, which provide extensive contextual information. These findings confirm that integrating various knowledge layers effectively boosts the capabilities of different backbones, including both smaller models like Vicuna-7b and larger models like GPT-4, thereby validating the effectiveness of our proposed approach.
V-F Query Proportion Addressed by Each Knowledge layer
To gain a deeper understanding of how the PolyRAG framework effectively impacts query outcomes, we conducted an analysis focusing on the proportion of queries addressed by different knowledge layers within the system. Specifically, we visualized the distribution of query resolutions across the various layers of our knowledge pyramid, including the raw text layer, the Knowledge Graph (KG) layer, and the Ontology layer. The results of this analysis are presented in Figure 6, which illustrates how each layer contributes to addressing user queries.
The visualization in Figure 6 reveals that a significant portion of queries, at least 28.5%, can be effectively addressed by the Ontology and KG layers. This finding highlights the crucial role that these structured knowledge layers play in enhancing the performance of the PolyRAG framework. The higher layers provide well-organized and semantically rich information that can directly satisfy user queries without the context for extensive processing. This efficiency not only improves response accuracy but also reduces computational overhead since the water-full framework will terminate when getting available contexts.
Moreover, the proportion of queries addressed by the Ontology and KG layers increases when more KG triplets (i.e., from top-5 to top-10) are included in the context provided to larger Language Model (LLM) backbones. Specifically, we observed that by including five additional KG triplets in the LLM prompt for answering questions, there is a notable increase of 7.6% (3%) in the number of queries resolved at the Ontology/KG layers by Vicuna-13b (7b). This demonstrates that enriching the context with more structured knowledge enables the LLMs to leverage this information more effectively, particularly in larger models capable of processing more extensive inputs.
This trend emphasizes the importance of integrating structured knowledge sources to enhance query processing efficiency and effectiveness. The Ontology and KG layers serve as valuable resources that can address a substantial portion of queries directly, reducing the reliance on raw text retrieval and complex language model reasoning. By augmenting the LLM prompts with additional KG triplets, we facilitate better comprehension and utilization of domain-specific knowledge by the language models, resulting in more accurate and contextually appropriate responses.
Figure 6: Query proportion addressed by Ontology, KG, Raw Text Layer on different LLM backbones.
VI Conclusion
This research outlines a preliminary investigation proposing a framework that integrates various knowledge representations to create a knowledge base and offers a multi-level query approach for applying the knowledge pyramid to domain-specific question answering. In this context, we addressed two significant challenges: the problem of prioritizing knowledge in a certain area and the presence of noise in retrieval situations. In order to overcome the limitations, we introduce Knowledge Pyramid multi-level retrieval framework as PolyRAG. It utilizes a sequential retrieval approach and prioritizes knowledge extraction, designed to tackle the challenges posed by the high demand for dense knowledge in domain-specific scenarios and the distractions caused by noisy contexts. The accuracy of PolyRAG has been validated through extensive tests done in AcadChall and R-FLUE-FiQA, surpassing earlier approaches and generating state-of-the-art results.